OpenStudy (izzy18970):

The approximate number of cell phones subscribers can be modeled by YC = 0.01(e^0.45x) x is the number of years since January 1980 and Yc is the number of subscribers in millions When did the number of subscribers reach 60 million?

OpenStudy (izzy18970):



OpenStudy (izzy18970):

@phi help pls

OpenStudy (phi):

60,000,000 = 0.01 exp(0.45 x) 6e7 = 1e-2 exp(0.45 x) (using scientific notation) 6e9 = exp(0.45 x) ln(6e9) = 0.45 x x = ln(6e9)/0.45 that gives x = 50.03 call it 50 years, so 1980+50 = 2030

OpenStudy (izzy18970):

ok gratzi

OpenStudy (izzy18970):

I have a 95 in math class and I got a 60 on a test grade. Test grades are 50% of your grade . Test average is 92